# WebGen-Bench: Evaluating LLMs on Generating Interactive and Functional Websites from Scratch

### v1

@@ Abstract
LLM-based agents have demonstrated great potential in generating and managing code within complex codebases. In this paper, we introduce WebGen-Bench, a novel benchmark designed to measure an LLM-based agent’s ability to create multi-file website codebases from scratch. It contains diverse instructions for website generation, created through the combined efforts of human annotators and GPT-4o. These instructions span three major categories and thirteen minor categories, encompassing nearly all important types of web applications. To assess the quality of the generated websites, we use GPT-4o to generate test cases targeting each functionality described in the instructions, and then manually filter, adjust, and organize them to ensure accuracy, resulting in 647 test cases. Each test case specifies an operation to be performed on the website and the expected result after the operation. To automate testing and improve reproducibility, we employ a powerful web-navigation agent to execute tests on the generated websites and determine whether the observed responses align with the expected results. We evaluate three high-performance code-agent frameworks—Bolt.diy, OpenHands, and Aider—using multiple proprietary and open-source LLMs as engines. The best-performing combination, Bolt.diy powered by DeepSeek-R1, achieves only 27.8% accuracy on the test cases, highlighting the challenging nature of our benchmark. Additionally, we construct WebGen-Instruct, a training set consisting of 6,667 website-generation instructions. Training Qwen2.5-Coder-32B-Instruct on Bolt.diy trajectories generated from a subset of this training set achieves an accuracy of 38.2%, surpassing the performance of the best proprietary model. We release our data-generation, training, and testing code, along with both the datasets and model weights at https://github.com/mnluzimu/WebGen-Bench.
@@ 1 Introduction
Recent developments in large language models (LLMs) have demonstrated increasingly strong performance. When paired with agent frameworks, they have become much more competent at solving challenging tasks such as fixing bugs in complex codebases and competing in coding competitions. Prior works have sought to quantify the software engineering abilities of these LLM-powered agents by testing them on curated GitHub issues [16, 39] and feature-patching requests [23]. These tasks involve advanced modifications to existing codebases and primarily target expert engineers.
Table 1: Comparison of WebGen-Bench with other repository-level software engineering benchmarks. indicates that the statistics for SWE-Bench Multimodal are median values, whereas the others are average values. The values for our benchmarks are gathered from the test results of Bolt.diy, OpenHands, and Aider using DeepSeek-V3. The values for the other benchmarks are taken from [16][23], and [39], respectively.
@@
On the other hand, there is a growing need for code agents to assist non-experts with little or no programming background in building applications tailored to their needs and expectations. For example, Bolt.new [^1] https://bolt.new and Lovable.dev [^2] https://lovable.dev are two projects that generate complete websites based on user requests and have become very popular among customers. This task poses significant challenges for LLM-based agents, as building a fully functional and customized web application from scratch tests a wide range of capabilities—including high-level planning, organizing complex multi-file codebases, and implementing nuanced user requirements. However, there is currently a lack of systematic and reliable evaluation methods for this task. The high demand for such applications, coupled with the value of assessing agent capabilities, highlights the need for a novel benchmark to evaluate the ability to generate websites from scratch based on natural language instructions.
@@
To this end, we introduce WebGen-Bench, the first benchmark to systematically evaluate LLM-based agents’ ability to construct websites that satisfy the functional and appearance requirements specified in user instructions. As shown in Table 1, unlike prior software-engineering benchmarks [16, 23, 39], which focus on fixing bugs or supplying patches to existing codebases, our benchmark requires models to build a complex codebase from scratch, assessing agents’ ability to plan, develop, and manage projects with multi-file structures. There are two critical challenges to address when creating the benchmark: (1) how to curate diverse instructions covering major web-application categories and (2) how to accurately evaluate the websites generated from scratch.
To tackle these problems, we introduce a systematic data curation and evaluation pipeline for assessing website-generation agents. Starting from 20 common categories identified across popular development platforms, we use GPT-4o to generate diverse instructions and test cases that cover both functionality and appearance, followed by manual verification. For evaluation, we leverage WebVoyager for automated functional testing and prompt GPT-4o to rate design aesthetics on a scale of 1 to 5. Using this framework, we benchmark Bolt.diy, OpenHands, and Aider, and find that Bolt.diy performs best. Further evaluation across models shows that DeepSeek-R1 achieves the highest functional success rate (27.8%), whereas Claude-3.5-Sonnet leads in appearance with an average score of 3.0, indicating substantial room for improvement.
@@
We also construct a training dataset named WebGen-Instruct, which contains 6,667 website-generation instructions. To avoid data contamination, we removed instructions that are semantically similar to those in WebGen-Bench by applying Jaccard-similarity filtering and Sentence-Transformers–based deduplication [28], as detailed in Appendix D. Fine-tuning Qwen2.5-Coder-32B-Instruct on Bolt.diy trajectories—generated from a subset of WebGen-Instruct by DeepSeek-V3 with rejection sampling raises its accuracy to 38.2%, a substantial improvement over its original 9.5% and even higher than the performance of DeepSeek-R1. We also fine-tune Qwen2.5-Coder-7B-Instruct and Qwen2.5-Coder-14B-Instruct on the same training data, and name the resulting family of website-generation models WebGen-LM.
@@ 2 Related Work
Software Engineering Benchmarks. Code generation has long been used as a means to evaluate the abilities of LLMs [12, 5, 4]. Previous works have collected coding problems from various sources, such as user queries [42], coding contests [15], model synthesis [44], and expert design [24, 5], to evaluate LLMs’ performance on single-file, function-level coding tasks. Recently, as stronger models have reached a plateau on these simpler benchmarks, new benchmarks such as SWE-bench [16, 39] and SWE-Lancer [23] have been constructed by collecting real-world code repositories and corresponding issue requests to test models’ ability to solve bugs and implement new functionalities. These benchmarks require models to identify and fix issues [16, 39, 2], perform code completions [18, 41], or provide functionality patches [23] within an existing multi-file codebase. Different from previous works, our benchmark focuses on creating web applications from scratch based on natural language instructions, requiring models to generate a complex, multi-file codebase, implement multiple functionality and appearance requirements, and make independent technical design decisions.
@@
LLM-based Code Agents and Pipelines. Various agent-based [33, 38, 1, 7, 6, 34] and pipeline-based [35, 29, 43] methods have been proposed to address software engineering problems such as code completion and GitHub issue resolution. While pipeline-based methods sometimes demonstrate strong performance on specific tasks with fixed pipelines [16], agent-based methods are generally more flexible. Code agent frameworks such as OpenHands [33] and SWE-agent [38] interact with executable environments to obtain feedback from the execution of generated code. To evaluate our benchmark, we selected three open-source code agents. Among them, OpenHands [33] and Aider [1] are general-purpose code agent frameworks that we adapted for our benchmark, while Bolt.diy [30] is a specialized framework for generating web applications. Prior works [26, 22, 20, 36, 21] have employed various post-training methods to improve the performance of open-source models. In this work, we also fine-tune open-source models with generated trajectories.
@@
Automatic Software User-testing. User-testing is a common method in software engineering to assess the functionality of software with high user-interaction requirements. However, human testing can be costly and introduce significant management complexities. Various works have employed agents to test websites [19], graphical user interfaces (GUIs) [8], and games [31, 9]. Among them, UXAgent [19] uses UI agents with pre-defined personas to simulate user experiences on websites. Our work also utilizes a web navigation UI agent to evaluate generated websites. Different from prior works, we define atomic test cases targeting functionality and appearance requirements, enabling the agent to perform operations and observe whether the website behaves as intended.
@@ 3.1 Instruction Curation
To ensure the diversity and practicality of the instructions, we first carefully browsed several platforms containing website development project listings, including Upwork [^3] https://www.upwork.com, Freelancer [^4] https://www.freelancer.com, and Proginn [^5] https://www.proginn.com. We identified twenty prevalent web application categories, as outlined in Table 8. To simulate numerous customized web applications, we employ a panel of forty computer science Ph.D students to conduct brainstorming sessions to determine various specific web applications belonging to these categories, as well as a brief and clear list of corresponding functionality and appearance requirements for each application. A customized application and its corresponding requirements are combined into a project description. We manually created 10152 project descriptions in total.
Website-Generation Instruction Curation.
@@
From the collected project descriptions, we use one-shot prompting with GPT-4o to generate the corresponding instructions. The prompt template is shown in Fig. 5 of Appendix C. Because the total number of generated instructions exceeds the practical limits of benchmarking code agents—which require substantial computational resources and long inference trajectories—we sample 2 to 8 representative examples from each category to preserve both coverage and diversity. This procedure produces a curated test set containing 101 instructions.
@@
Next, we decontaminate the remaining instructions by first filtering those with a 5-gram Jaccard similarity score exceeding 0.6 relative to any testing instruction. We then perform semantic deduplication by computing cosine similarity between sentence embeddings [28] of the remaining instructions and the testing set. This process produces a training set of 6,667 website-generation instructions, which we name WebGen-Instruct. Details of the decontamination process are provided in Appendix D.
Test Set Adjustment and Validation.
@@
We refine and validate the selected test instructions to ensure they exclude unreasonable designs and specific technical details. We intentionally omit technical design specifications because our dataset aims to evaluate code agents in scenarios where they receive instructions from non-expert users. The agents should autonomously determine the optimal technical approach. Including tool-specific hints in the instructions would compromise this objective.
Technical Classification of the Testing Set.
@@
Given the limited number of testing instructions per application category, analyzing categorical statistics based on the original 20 application categories would be confusing. To enable higher-level analysis, we reorganize the 101 testing instructions into three broader technical categories (see Tab. 2): 1. Content Presentation: Static page generation (e.g., corporate/portfolio sites), dynamic rendering (e.g., blogs/news feeds), data visualization (e.g., dashboards), and immersive media displays (e.g., 360° product views). 2. User Interaction: Form systems, authentication flows, real-time collaboration tools, e-commerce transactions, and AI-enhanced features (e.g., chatbots). 3. Data Management: CRUD operations for content administration, third-party API integrations (e.g., payment/social platforms), analytical processing of user behavior data, and file operations (e.g., cloud synchronization, bulk exports).
@@ 3.2 Test Case Construction and Evaluation
Test Case Construction.
Each test case consists of an operation verifying a specific functionality or appearance requirement, paired with its expected outcome. We first generate draft test cases using GPT-4o with the prompt shown in Fig. 10. Two computer science Ph.D. students then independently review and refine these test cases. After comparing their adjustments, we resolved discrepancies through discussion, yielding a final set of 647 test cases (4–11 per instruction). This manual validation process guarantees strict alignment between test cases and instructions, ensuring: (1) all instruction requirements are covered by test cases, and (2) each test case corresponds to an instruction requirement. This approach ensures comprehensive evaluation while preserving implementation flexibility for the tested agents.
UI Agent-based Evaluation.
With instructions and test cases prepared, we must determine how to effectively evaluate the generated websites. Manual testing by human evaluators is costly and time-consuming, as completing a test case takes at least 60 seconds, and finishing all 647 test cases would require approximately 10.8 hours at an estimated cost of \$377.8 [32]. This slow, labor-intensive process would hinder rapid iteration during framework development, preventing researchers from obtaining timely feedback when refining website-generation systems.
To improve testing efficiency, we automate test case evaluation. Inspired by [19], which employs persona-based agents for web usability testing, we utilize WebVoyager [11], a robust web navigation UI agent, to execute test operations and verify outcomes. We structure each test case’s operation and expected outcome into a standardized prompt (Fig. 12), which directs the agent to simulate user interactions, analyze action trajectories and screenshots, and return YES, NO, or PARTIAL assessments based on requirement fulfillment. The process is shown on the right side of Fig. 1 (b). When the agent reaches its interaction limit, we trigger a decision prompt, inducing the agent to make a final decision (Fig. 11). Considering the cost induced by multiple interactions with the website in evaluating each test case, we employ Qwen2.5-VL-32B-Instruct, an efficient open-source vision-language model that balances performance and cost-effectiveness, as the agent’s engine.

### v2

@@ Abstract
LLM-based agents have demonstrated great potential in generating and managing code within complex codebases. In this paper, we introduce WebGen-Bench, a novel benchmark designed to measure an LLM-based agent’s ability to create multi-file website codebases from scratch. It contains diverse instructions for website generation, created through the combined efforts of human annotators and GPT-4o. These instructions span three major categories and thirteen minor categories, encompassing nearly all important types of web applications. To assess the quality of the generated websites, we use GPT-4o to generate test cases targeting each functionality described in the instructions, and then manually filter, adjust, and organize them to ensure accuracy, resulting in 647 test cases. Each test case specifies an operation to be performed on the website and the expected result after the operation. To automate testing and improve reproducibility, we employ a powerful web-navigation agent to execute tests on the generated websites and determine whether the observed responses align with the expected results. We evaluate three high-performance code-agent frameworks—Bolt.diy, OpenHands, and Aider—using multiple proprietary and open-source LLMs as engines. The best-performing combination, Bolt.diy powered by DeepSeek-R1, achieves only 27.8% accuracy on the test cases, highlighting the challenging nature of our benchmark. Additionally, we construct WebGen-Instruct, a training set consisting of 6,667 website-generation instructions. Training Qwen2.5-Coder-32B-Instruct on Bolt.diy trajectories generated from a subset of this training set achieves an accuracy of 38.2%, surpassing the performance of the best proprietary model. We release our data-generation, training, and testing code, along with both the datasets and model weights at https://github.com/mnluzimu/WebGen-Bench, to facilitate future research.
Equal contribution Corresponding author §Project lead
@@ 1 Introduction
Recent developments in large language models (LLMs) have demonstrated increasingly strong performance. When paired with agent frameworks, they have become much more competent at solving challenging tasks such as fixing bugs in complex codebases and competing in coding competitions. Prior works have sought to quantify the software engineering abilities of these LLM-powered agents by testing them on curated GitHub issues (jimenez2023swe, ; yang2024swe, ) and feature-patching requests (miserendino2025swe, ). These tasks involve advanced modifications to existing codebases and primarily target expert engineers.
Table 1: Comparison of WebGen-Bench with other repository-level software engineering benchmarks. indicates that the statistics for SWE-Bench Multimodal are median values, whereas the others are average values. The values for our benchmarks are gathered from the test results of Bolt.diy, OpenHands, and Aider using DeepSeek-V3. The values for the other benchmarks are taken from jimenez2023swe miserendino2025swe , and yang2024swe , respectively.
@@
On the other hand, there is a growing need for code agents to assist non-experts with little or no programming background in building applications tailored to their needs and expectations. For example, Bolt.new [^1] https://bolt.new and Lovable.dev [^2] https://lovable.dev are two projects that generate complete websites based on user requests and have become very popular among customers. This task poses significant challenges for LLM-based agents, as building a fully functional and customized web application from scratch tests a wide range of capabilities—including high-level planning, organizing complex multi-file codebases, and implementing nuanced user requirements in both functionality and design. However, there is currently a lack of systematic and reliable evaluation methods for this task. The high demand for such applications, coupled with the value of assessing agent capabilities, highlights the need for a novel benchmark to evaluate the ability to generate websites from scratch based on natural language instructions.
@@
To this end, we introduce WebGen-Bench, the first benchmark to systematically evaluate LLM-based agents’ ability to construct websites that satisfy the functional and appearance requirements specified in user instructions. As shown in Table 1, unlike prior software-engineering benchmarks jimenez2023swe ; miserendino2025swe ; yang2024swe , which focus on fixing bugs or supplying patches to existing codebases, our benchmark requires models to build a complex codebase from scratch, assessing agents’ ability to plan, develop, and manage projects with multi-file structures. There are two critical challenges to address when creating the benchmark: (1) how to curate diverse instructions covering major web-application categories and (2) how to accurately evaluate the websites generated from scratch. WebGen-Bench tackles these problems by providing a systematic data-curation process and a reliable evaluation pipeline that thoroughly assesses the agents’ website-generation abilities.
To construct the benchmark dataset, we initially examined prominent platforms listing website-development projects, such as Upwork [^3] https://www.upwork.com/, Freelancer [^4] https://www.freelancer.com/, and Proginn [^5] https://www.proginn.com. Through systematic analysis and discussions among the authors, we identified twenty common categories of website-development requirements. Subsequently, a panel of forty computer-science Ph.D. students conducted extensive brainstorming sessions to determine numerous specific web applications for each category, along with brief, clear functionality and appearance requirements for each application. Then, we employed GPT-4o to convert these web applications and their corresponding requirements into diverse website-generation instructions. Next, we prompted GPT-4o to generate test cases for each instruction, targeting both functional and appearance requirements. Each test case consists of an operation to be performed on the website and the expected outcome. Finally, we manually reviewed each test case, verifying that it is reasonable and comprehensively covers all requirements specified in the corresponding instruction.
Using human testers for evaluation is slow, costly, and presents management challenges. Inspired by lu2025uxagent , which employs agents with specified personas for automated testing, we utilize a strong web navigation agent, WebVoyager he2024webvoyager , to perform operations and assess outcomes. Furthermore, we also evaluates the harmony and aesthetic quality of the website design by designing a prompt that asks GPT-4o to grade the appearance of the generated websites on a scale from 1 to 5. This provides a comprehensive assessment of overall website generation quality. The testing process is demonstrated in Fig. 1 (b). With this testing framework, we evaluate the performance of three strong open-source code agents — Bolt.diy, OpenHands, and Aider — using DeepSeek-V3 as the engine LLM. Among them, Bolt.diy achieves the best performance. We further evaluate Bolt.diy across a diverse set of proprietary and open-source models. DeepSeek-R1 achieves the best performance, attaining an accuracy of only 27.8% on the test cases, while Claude-3.5-Sonnet achieves the highest average appearance score of 3.0. The low accuracies and appearance scores demonstrate that the performance of current LLM-based agents remains far from saturated on our benchmark.
@@
We also construct a training dataset named WebGen-Instruct, which contains 6,667 website-generation instructions. To avoid data contamination, we removed instructions that are semantically similar to those in WebGen-Bench by applying Jaccard-similarity filtering and Sentence-Transformers–based deduplication (reimers-2019-sentence-bert, ), as detailed in Appendix D. Fine-tuning Qwen2.5-Coder-32B-Instruct on Bolt.diy trajectories—generated from a subset of WebGen-Instruct by DeepSeek-V3 with rejection sampling raises its accuracy to 38.2%, a substantial improvement over its original 9.5% and even higher than the performance of DeepSeek-R1. We also fine-tune Qwen2.5-Coder-7B-Instruct and Qwen2.5-Coder-14B-Instruct on the same training data, and name the resulting family of website-generation models WebGen-LM.
@@ 2 Related Work
Software Engineering Benchmarks. Code generation has long been used as a means to evaluate the abilities of LLMs (hendrycks2021measuring, ; chen2021evaluating, ; austin2021program, ). Previous works have collected coding problems from various sources, such as user queries (zhang2024naturalcodebench, ), coding contests (jain2024livecodebench, ), model synthesis (zhuo2024bigcodebench, ), and expert design (muennighoff2023octopack, ; chen2021evaluating, ), to evaluate LLMs’ performance on single-file, function-level coding tasks. Recently, as stronger models have reached a plateau on these simpler benchmarks, new benchmarks such as SWE-bench (jimenez2023swe, ; yang2024swe, ) and SWE-Lancer (miserendino2025swe, ) have been constructed by collecting real-world code repositories and corresponding issue requests to test models’ ability to solve bugs and implement new functionalities. These benchmarks require models to identify and fix issues (jimenez2023swe, ; yang2024swe, ; aleithan2024swe, ), perform code completions (liu2023repobench, ; zhang2023repocoder, ), or provide functionality patches (miserendino2025swe, ) within an existing multi-file codebase. Different from previous works, our benchmark focuses on creating web applications from scratch based on natural language instructions, requiring models to generate a complex, multi-file codebase, implement multiple functionality and appearance requirements, and make independent technical design decisions.
@@
LLM-based Code Agents and Pipelines. Various agent-based (wang2024openhands, ; yang2024sweagent, ; aiderai2024aider, ; Cursor2024cursor, ; GitHub2024copilot, ; Wu2024devin, ) and pipeline-based (xia2024agentless, ; ruan2024specrover, ; zhang2024autocoderover, ) methods have been proposed to address software engineering problems such as code completion and GitHub issue resolution. While pipeline-based methods sometimes demonstrate strong performance on specific tasks with fixed pipelines (jimenez2023swe, ), agent-based methods are generally more flexible and versatile. Code agent frameworks such as OpenHands (wang2024openhands, ) and SWE-agent (yang2024sweagent, ) interact with executable environments to obtain feedback from the execution of generated code. To evaluate our benchmark, we selected three open-source code agents. Among them, OpenHands (wang2024openhands, ) and Aider (aiderai2024aider, ) are general-purpose code agent frameworks that we adapted for our benchmark, while Bolt.diy (stackblitzlabs2024bolt, ) is a specialized framework for generating web applications. Prior works (pan2024training, ; ma2025sorft, ; ma2024lingma, ; xie2025swe, ; ma2024repository, ) have employed various post-training methods to improve the performance of code agents using open-source models. In this work, we also fine-tune open-source models with trajectories generated from web-generating instructions.
@@
Automatic Software User-testing. User-testing is a common method in software engineering to assess the functionality of software with high user-interaction requirements. However, human testing can be costly and introduce significant management complexities. Various works have employed agents to test websites lu2025uxagent , graphical user interfaces (GUIs) (9196452, ), and games (10.1145/3290607.3313039, ; 9440143, ). Among them, UXAgent (lu2025uxagent, ) uses UI agents with pre-defined personas to simulate user experiences on websites. Our work also utilizes a web navigation UI agent to evaluate generated websites. Different from prior works, we define atomic test cases targeting functionality and appearance requirements, enabling the agent to perform operations and observe whether the website behaves as intended.
@@ 3.1 Instruction Curation
Web Development Project Descriptions Collection. To ensure the diversity and practicality of the instructions, we first carefully browsed several platforms containing website development project listings, including Upwork [^6] https://www.upwork.com, Freelancer [^7] https://www.freelancer.com, and Proginn [^8] https://www.proginn.com. After systematic analysis and discussions among the authors, we identified twenty prevalent web application categories, as outlined in Table 7. To simulate numerous customized web applications, we employ a panel of forty computer science Ph.D students to conduct brainstorming sessions to determine various specific web applications belonging to these categories, as well as a brief and clear list of corresponding functionality and appearance requirements for each application. A customized application and its corresponding requirements are combined into a project description. We manually created 10152 project descriptions in total.
Web Development Project Descriptions Collection. To ensure the diversity and practicality of the instructions, we first carefully browsed several platforms containing website development project listings, including Upwork [^9] https://www.upwork.com, Freelancer [^10] https://www.freelancer.com, and Proginn [^11] https://www.proginn.com. After systematic analysis and discussions among the authors, we identified twenty prevalent web application categories, as outlined in Table 7. To simulate numerous customized web applications, we employ a panel of forty computer science Ph.D students to conduct brainstorming sessions to determine various specific web applications belonging to these categories, as well as a brief and clear list of corresponding functionality and appearance requirements for each application. A customized application and its corresponding requirements are combined into a project description. We manually created 10152 project descriptions in total.
@@
Website-Generation Instruction Curation. From the collected project descriptions, we use one-shot prompting with GPT-4o to generate the corresponding instructions. The prompt template is shown in Fig. 4 in Appendix C. Because the total number of generated instructions exceeds the practical limits of benchmarking code agents—which require substantial computational resources and long inference trajectories—we sample two to eight representative examples from each category to preserve both coverage and diversity. This procedure produces a curated test set containing 101 instructions.
@@
Next, we deduplicate the remaining instructions by first filtering those with a 5-gram Jaccard similarity score exceeding 0.6 relative to any testing instruction. We then perform semantic deduplication by computing cosine similarity between sentence embeddings (reimers-2019-sentence-bert, ) of the remaining instructions and the testing set. This decontamination process produces a training set of 6,667 website-generation instructions, which we name WebGen-Instruct. Details of the decontamination process are provided in Appendix D.
@@
Test Set Adjustment and Validation. We refine and validate the selected test instructions to ensure they exclude unreasonable designs and specific technical implementation details. We intentionally omit technical design specifications because our dataset aims to evaluate code agents in scenarios where they receive instructions from non-expert users. The agents should autonomously determine the optimal technical approach. Including tool-specific hints in the instructions would compromise this objective.
@@
Technical Classification of the Testing Set. Given the limited number of testing instructions per application category, analyzing categorical statistics based on the original 20 application categories would be confusing. To enable higher-level analysis, we reorganize the 101 testing instructions into three broader technical categories (see Tab. 2:) 1. Content Presentation: Static page generation (e.g., corporate/portfolio sites), dynamic rendering (e.g., blogs/news feeds), data visualization (e.g., dashboards), and immersive media displays (e.g., 360° product views). 2. User Interaction: Form systems, authentication flows, real-time collaboration tools, e-commerce transactions, and AI-enhanced features (e.g., chatbots). 3. Data Management: CRUD operations for content administration, third-party API integrations (e.g., payment/social platforms), analytical processing of user behavior data, and file operations (e.g., cloud synchronization, bulk exports).
@@ 3.2 Test Case Construction and Evaluation
Test Case Construction. Each test case consists of an operation verifying a specific functionality or appearance requirement, paired with its expected outcome. We first generate draft test cases using GPT-4o with the prompt shown in Figure 8. Two computer science Ph.D. students then independently review and refine these test cases. After comparing their adjustments, we resolved discrepancies through discussion, yielding a final set of 647 test cases (4–11 per instruction). This manual validation process guarantees strict alignment between test cases and instructions, ensuring: (1) all instruction requirements are covered by test cases, and (2) each test case corresponds to an instruction requirement. This approach ensures comprehensive evaluation while preserving implementation flexibility for the tested agents.
UI Agent-based Evaluation. With instructions and test cases prepared, we must determine how to effectively evaluate the generated websites. Manual testing by human evaluators is costly and time-consuming, as completing a test case takes at least 60 seconds, and finishing all 647 test cases would require approximately 10.8 hours at an estimated cost of \$377.8 (uswage, ). This slow, labor-intensive process would hinder rapid iteration during framework development, preventing researchers from obtaining timely feedback when refining website-generation systems.
To improve testing efficiency, we automate test case evaluation. Inspired by lu2025uxagent , which employs persona-based agents for web usability testing, we utilize WebVoyager he2024webvoyager , a robust web navigation UI agent, to execute test operations and verify outcomes. We structure each test case’s operation and expected outcome into a standardized prompt (Fig. 10), which directs the agent to simulate user interactions, analyze action trajectories and screenshots, and return YES, NO, or PARTIAL assessments based on requirement fulfillment. The process is shown on the right side of Fig. 1 (b). When the agent reaches its interaction limit without completing evaluation, we trigger a decision prompt, inducing the agent to make a final decision (Fig. 9). Considering the cost induced by multiple interactions with the website in evaluating each test case, we employ Qwen2.5-VL-32B-Instruct, an efficient open-source vision-language model that balances performance and cost-effectiveness, as the agent’s engine.